OpenStudy (anonymous):

a) given the quadratic equation hx^2+kx+6=0 where h and k are constant. i) if the equation has two equal roots, express h in terms of k. ii) if h=6 and the quadratic equation has two different roots, find the range of values of k b) given the straight line y=mx+3 does not intercept the curve y=2x^2 +x+5. find the range of values of m.

OpenStudy (blockcolder):

If it has 2 equal roots, then it means that discriminant=0. If it has 2 distinct roots, then discriminant>0.

OpenStudy (blockcolder):

Want me to elaborate further? :)

OpenStudy (anonymous):

@blockcolder yes sure.

OpenStudy (blockcolder):

\[\large \sqrt{k^2-24h}=0 \Rightarrow k^2=24h \Rightarrow h=\frac{k^2}{24}\\ \large \sqrt{k^2-24h}>0 \Rightarrow \sqrt{k^2- 144}>0 \Rightarrow k^2>144 \Rightarrow k<-12 \text{ or } k>12\].

OpenStudy (anonymous):

Cant get it..huhu how u get the 24 h?

OpenStudy (blockcolder):

\[\text{Discriminant }=\sqrt{b^2-4ac}\] In this case, a=h, b=k, and c=6. :)

OpenStudy (anonymous):

@Daenio @Callisto @Kreshnik @Mani_Jha CAN HELP ME, PLEASE?

OpenStudy (saifoo.khan):

Where are you stuck now? @gF

OpenStudy (anonymous):

from the start...huhu

OpenStudy (mani_jha):

As blockcolder already said, equal roots mean that the value of b^2-4ac is 0. a,b, are the coefficients of x^2, x terms and c is the constant one. In the given equation, the coefficient of x^2 is h, of x is k, and the constant term is 6. So a=h b=k and c =6 Now b^2-4ac=0 or, k^2-4*6*h=0 Now ok for the first part?

OpenStudy (anonymous):

yes..then?

OpenStudy (mani_jha):

\[k ^{2}=24h\] \[h=k ^{2}/24\] For the next part, if the roots are different, b^2-4ac is not zero but is greater than zero(which means it is positive). k^2-24h>0 Substitute the given value of h=6